A teacher is 15 years older than her student. In 15 years, the ratio of their ages will be 3:2. Find their present ages.
Answer by drk(1908)

(i) T = 15 + S
from above we get
(T+15)/(s+15) = 3/2
by substitution we get
(15+S +15)/(S+15) = 3/2
cross multiply to get
2(S+30) = 3(S+15)
2S + 60 = 3S + 45
60 = S + 45
So, s = 15
and T = 30